During the combustion of 2.8g C, the temperature of 1200g of water in a calorimeter rose from 18.4 degrees Celsius to 55.7 degrees Celsius. Determine the molar heat for this reaction
1
Expert's answer
2019-03-29T03:33:12-0400

The formula of quantity of heat, q, is equal to the mass of substance, m, multiplied with the specific heat and the change in temperature, ΔT .

C(H2O) = 4.186 J/g·°C

q = C*m*ΔT = 4.186 J/g·°C*1200g*(55.7°C – 18.4°C) = 187365.36 J = 187.365 kJ

n(C) = m/M = 2.8g/12g/mol = 0.233 mol

Molar heat = 187.365kJ/0.233mol = 804.142 kJ/mol